OpenStudy (lena772):

Sandra uses her office fax machine to send fax at the rate of $0.10 per page. She decides to rent a fax machine for $80 a year. The cost of sending a fax using the rented machine is $0.06 per page. Part A: Write an inequality that can be used to calculate the number of pages that Sandra should fax in a year so that the amount she pays for the rented machine is less than the office machine. Define the variable used. (5 points) Part B: How many pages should Sandra fax in a year to justify renting the fax machine? Show your work. (5 points)

ganeshie8 (ganeshie8):

Her office fax machine seems to be using so much power/watever the fax machine eats to send the fax : $0.10 per page !

ganeshie8 (ganeshie8):

So she decides to rent a smart fax machine for an year, which can send the fax at a very low price : just $0.06 per page

ganeshie8 (ganeshie8):

you get this far eh ?

ganeshie8 (ganeshie8):

But the new fax machine doesnt come freely, she has to pay rent of $80 per year

ganeshie8 (ganeshie8):

say she prints \(x\) pages in a year, so the cost for printing these pages using new RENTED fax machine = ?

OpenStudy (lena772):

80+0.06x

ganeshie8 (ganeshie8):

Perfect ! so, how much it would have costed, if she kept using the Office dumb fax machine oly ?

ganeshie8 (ganeshie8):

cost for prining \(x\) pages using the existing office fax machine : $0.1x

OpenStudy (lena772):

0.10x

ganeshie8 (ganeshie8):

to break even, we need the cost of RENTED machine to be LESS than this guy ok

ganeshie8 (ganeshie8):

so, set up an inequality

ganeshie8 (ganeshie8):

can u ? :)

OpenStudy (lena772):

80+0.06x=0.10x?

ganeshie8 (ganeshie8):

*inequality girly

ganeshie8 (ganeshie8):

\(80+0.06x \color{red}{<} 0.10x\)

ganeshie8 (ganeshie8):

cost wid NEW rented fax machine must have to be less than the cost for faxing wid the OLD office machine. then oly she can justify to her boss renting the new machine

OpenStudy (lena772):

80+0.06x < 0.10x 80<0.04x

ganeshie8 (ganeshie8):

Yes ! move to part B, and simply solve the inequality for x

OpenStudy (lena772):

3000<x

ganeshie8 (ganeshie8):

careful

ganeshie8 (ganeshie8):

80<0.04x 8000 < 4x 8000/4 < x 2000 < x

OpenStudy (anonymous):

so part b is 2000 pages? @ganeshie8

OpenStudy (anonymous):

and where did you get 8000

ganeshie8 (ganeshie8):

yes 2000 or greater

ganeshie8 (ganeshie8):

80<0.04x multiply both sides by 100

ganeshie8 (ganeshie8):

you wud get 8000 on left side, 4x on right side

OpenStudy (anonymous):

could i show you another question, im already done i just wanna make sure its right.

ganeshie8 (ganeshie8):

sure :)

OpenStudy (anonymous):

This is the question: Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points) Part B: How many liters of apple juice are present in the jar? Show your work. (5 points)

OpenStudy (anonymous):

Part A: The total volume be x liters There's 75% apple juice, (75/100)*x 0.75x total volume=0.75x+1 x=0.75x+1 Part B: 0.25x=1 x=4 there's 0.75x liters apple juice present in the container 0.75*x

ganeshie8 (ganeshie8):

Part A is good, let me adjust Part B a bit

ganeshie8 (ganeshie8):

Part B: 0.25x=1 x=4 there's 0.75x liters apple juice present in the container 0.75*x = 0.75*4 = 3 So, there is 3 liters of apple juice present in the container

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

u wlc !